Can someone please help me solve these 3 problems....I'm still lost. Thanks:)
1) For a home that has a roof that is 7m from peak to eave and the attic space is 3m high at the peak, how long does each of the pieces of insulation need to be? Round to the nearest tenth.
2) Find the distance between each pair of roots.
(-3, 0) and (4, 0)

3) Use the Pythagorean theorem to determine the length of each line segment. Where appropriate, round to the nearest hundredth.
(3,2) (1,-2)

Answer by stanbon(75887)

1) For a home that has a roof that is 7m from peak to eave and the attic space is 3m high at the peak, how long does each of the pieces of insulation need to be? Round to the nearest tenth.
--------
Draw the picture. I assume you want to put the insulation horizontally on
the floor of the attic.
The horizontal distance = sqrt(7^2-3^2)= sqrt40 = 2sqrt10 ft.
2sqrt10= 6.32 ft
----------------------
2) Find the distance between each pair of roots.
(-3, 0) and (4, 0)
d=sqrt[(0-0)^2+(4--3)^2 = sqrt[49]=7
-------------------------
3) Use the Pythagorean theorem to determine the length of each line segment. Where appropriate, round to the nearest hundredth.
(3,2) (1,-2)
d=sqrt[(-2-2)^2 + (1-3)^2] = sqrt[16+4]=sqrt20=2sqrt5
===================
Cheers,
stan H.